# Constraining $\boldsymbol{|V_{cs}|}$ and physics beyond the Standard Model from exclusive (semi)leptonic charm decays

### v1

@@ A.3 Effective theory relations for baryon form factors
The $\Lambda_{c}\to\Lambda$ tensor form factors are available from a single lattice QCD analysis Meinel:2016dqj that provides the (axial)vector and (pseudo)scalar form factors, i.e., all form factors needed to produce theory predictions within the SM. For the purpose of our analysis as discussed in section 2.4, knowledge of the tensor form factors is also required. Information on the latter is presently not available from lattice QCD analyses. In the absence of such lattice QCD information, we rely on other constraints.
@@
Although the values of $\xi,\xi_{1}$, and $\xi_{2}$ can be approximated, a proper estimation of the correlation between the different form factors would require dedicated analyses. Here, we instead follow the recipe proposed in Ref. Amhis:2022vcd and restrict ourselves to the following set of relations, valid both at $q^{2}=q^{2}_{\mathrm{max}}$ (HQL) and $q^{2}=0$ (LEL):

### v2

@@ A.3 Effective theory relations for baryon form factors
The $\Lambda_{c}\to\Lambda$ tensor form factors are available from a single lattice QCD analysis Meinel:2016dqj that provides the (axial)vector and (pseudo)scalar form factors, i.e., all form factors needed to produce theory predictions within the SM. For the purpose of our analysis, as discussed in section 2.4, knowledge of the tensor form factors is also required. Information on the latter is presently not available from lattice QCD analyses. In the absence of such lattice QCD information, we rely on other constraints.
@@
Although the values of $\xi,\xi_{1}$ and $\xi_{2}$ can be approximated, a proper estimation of the correlation between the different form factors would require dedicated analyses. Here, we instead follow the recipe proposed in Ref. Amhis:2022vcd and restrict ourselves to the following set of relations, valid both at $q^{2}=q^{2}_{\mathrm{max}}$ (HQL) and $q^{2}=0$ (LEL):